The area of a typical eardrum is about 5.00 x 10^-
Answer #1

given

A = 5*10^-5 m^2

I = 1 W/m^2

a) we know, Intensity = power/Area

==> Power = Intensity*Area

= 1*5*10^-5

= 5*10^-5 W

b) Energy transefred, E = Power*time taken

= 5*10^-5*3.05*60

= 9.15*10^-3 J